OpenStudy (anonymous):

Which of the following expressions are equivalent? Justify your reasoning. A.4√x3 B.1x^−1 C.10√x 5•x4•x 2 x D. 1/3•x1/3•x1/3

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

@mathstudent55

OpenStudy (ali2x2):

is this the full question?

OpenStudy (anonymous):

@ali2x2 yes.

OpenStudy (ali2x2):

ok

OpenStudy (ali2x2):

A = 4√x3 = 4√x^2 x = 4 x √x B = 1/x^(-1) = x C = 10 √x^11 = √x^10 x = 10 x^5√x D = x^(1/3+1/3+1/3) = x^(3/3) = x B and D are equivalent

OpenStudy (anonymous):

@ali2x2 thanks!

OpenStudy (ali2x2):

mhm ;) just remember to close this question for future questions you need to ask. :)